OpenStudy (anonymous):

f(x)=3x-2 and g(x)= x-6. find (f*g)(x). give your answer in expanded form. (f*g)(x)=

OpenStudy (anonymous):

f*g(x) = f(g(x)) g(x) = x-6 f(x) = 3x - 2 f(x-6) = 3(x-6)-2 = 3x - 20